FIRST ASSIGNMENT OF SEVERAL TORSIONAL BANDS OF CH_{2}DOH

M. SMIRNOV, \underline{L. H. COUDERT}, LISA, CNRS/Universités Paris 12 et 7, 61 Avenue du Général de Gaulle, 94010 Créteil, France; S. BLASCO, D. LAUVERGNAT, Laboratoire de Chimie Physique, Unité mixte du C.N.R.S. et de l’Université Paris-Sud, Bât. 490, Université Paris-Sud, 91405 Orsay Cedex, France; and S. KLEE, Physikalisch-Chemisches Institut, Justus-Liebig-Universität Gießen, 35392 Gießen, Germany.

Although the torsional spectrum of the isotopic species of methanol with a symmetrical CH_{3} or CD_{3} methyl group has been studied in great details, almost no investigations have been carried out on partially deuterated species of methanol, like CH_{2}DOH or CD_{2}HOH, with an asymmetrical CH_{2}D or CD_{2}H methyl group. This stems from the fact that these partially deuterated species of methanol require a complicated theoretical treatment^{a} in order to compute their rotation-torsion energy levels and that they display a dense far infrared torsional spectrum^{b} difficult to assign. With a view toward understanding this spectrum, a theoretical calculation of the rotation-torsion energy levels of CH_{2}DOH has been undertaken aided by ab initio calculations. This approach accounts for the complicated torsion-rotation interaction displayed by this molecule and for the fact that, due to the low symmetry of the CH_{2}D methyl group, the inertia tensor is strongly dependent on the torsional angle of rotation. In the theoretical approach, this angle is treated as an active coordinate^{c} and matrix elements of the kinetic energy part of the Hamiltonian, involving the inertia tensor, are calculated using Gaussian quadrature.^{d} Rotation-torsion energy levels are obtained using a potential energy function retrieved through ab initio calculations. In the paper, a calculated rotation-torsion spectrum will be presented and compared to the experimental one, recorded in Gießen. It will be shown that this allows us to assign about 50 torsional bands.

aQuade and Lin, J. Chem. Phys. 38, 540 (1963).

bMukhopadhyay, Perry, Butler, Herbst, and De Lucia, 57th International Symposium on Molecular Spectroscopy, paper RH06 (2002) and Mukhopadhyay, Perry, Lock, and Klee, 57th International Symposium on Molecular Spectroscopy, paper RH07 (2002).

cLauvergnat and Nauts, J. Chem. Phys. 116, 8560 (2002).

dLight and Bačić, J. Chem. Phys. 87, 408 (1987).